suppose 3 companies x y z produce tvs x pproduces twice as many as x while y and z produce same number as.it is known that 2%of x,2%of y,4%of z are defected all the tvs are produced are put into 1 shop and then 1 tv is choosen at random what is the probability that the tv is defeceted.suppppose a tv choosen is defective what is the probability that this tv is produced by company x.

1
Expert's answer
2021-11-08T09:02:02-0500

Let "X" denote the event "tv is produced by company x."

Let "Y" denote the event "tv is produced by company y."

Let "Z" denote the event "tv is produced by company z."

Let "D" denote the event "tv is defected".

Given

"P(X)=0.5, P(Y)=P(Z)=0.25"

"P(D|X)=0.02, P(D|Y)=0.02,P(D|Z)=0.04"

a) Theorem of total probability


"P(D)=P(X)P(D|X)+P(Y)P(D|Y)"

"+P(Z)P(Z|D)"

"=0.5(0.02)+0.25(0.02)+0.25(0.04)=0.025"

The probability that the tv is defected is "0.025."


ii) By the Bayes' Rule


"P(X|D)"

"=\dfrac{P(X)P(D|X)}{P(X)P(D|X)+P(Y)P(D|Y)+P(Z)P(Z|D)}"

"=\dfrac{0.5(0.02)}{0.5(0.02)+0.25(0.02)+0.25(0.04)}"

"=0.4"

The probability that this tv is produced by company x is "0.4."